# HPVM-HDC: A Heterogeneous Programming System for Accelerating Hyperdimensional Computing

### v2

@@ 5.3. Impact of HDC optimizations
For this evaluation, we set the encoding dimension to 10240 and derived the trained class hypervectors using cosine similarity in a single pass over the training dataset offline. Figure 7 plots the relative speed-up versus inference accuracy for different optimization settings described in Table 3. These configurations are not meant to be exhaustive but merely illustrative of points in the various stages (e.g. encoding and/ or similarity) in an HDC application which can be approximated. The baseline (I) achieves an inference accuracy of 0.82. While all optimization settings achieve a significant performance improvement, the inference accuracy varies drastically ranging from a low of 0.2 to a high of 0.89. The points in green (III, VII, VIII) result in improved accuracy over the baseline, as well as faster performance. When using the binarization approximation, i.e., projecting the floating point values in the class hypervectors to {+1, -1}, which are represented using 1-bit elements, and using Hamming distance similarity measure, we observe a negligible change in accuracy. This aligns with observations in (Imani et al., 2019c) regarding the effects of binarized learning on accuracy. Additionally, this reduces the initial data movement necessary as the class hypervectors are represented using fewer bits in total. Reduction perforation adds upon the performance improvements achieved from automatic binarization. However, the downstream effect on accuracy varies significantly depending on which operation is being perforated. Configurations (VII, VIII, X) perforate the inference’s similarity computation and result in at most an approximately 10% drop in accuracy with respect to the baseline in X and no drop in the other two. However, when perforating the encoding stages (V, VI, IX), the inference accuracy drops to 25-35%. This implies that for HD classification, the similarity computation is much more amenable to approximation than the encoding stage.
@@ 5.4. Programmability
Lines of Code: The high-level HDC operations in HDC++ also enable more concise code. We compare the lines of code required to implement the applications with their respective baselines. Note that for each target architecture, there is a different baseline version, often implemented in a different programming language. Table  4 describes the total lines of code reduction in HDC++ versus the baselines for the applications. HDC++ provides a 1.6x geomean reduction in total lines of code (by combining the lines of codes across all baseline target implementations) for HDC applications while remaining retargetable. For the CPU baselines (except for HyperOMS where no baseline exists) which are implemented in Python, the HDC++ application codes are larger. This is a consequence of HDC++ being embedded in C++ which is inherently more verbose than Python. For the CUDA baselines however, HDC++ code is more compact; HDC++ abstracts away any data-movement and allocation between the device and host-memory. Further, the high-level HDC++ primitives compactly represent the corresponding manually written kernel functions in the GPU baselines, resulting in fewer lines of code. Note that not all baselines provide implementations for all targets. In contrast, HPVM-HDC can compile a single HDC++ application to multiple targets.
Approximation Configuration: The HDC++ programming interface and HPVM-HDC compiler make it easy to fine-tune HDC applications for performance and accuracy because the approximation configuration can be separately specified with almost no changes to the application source code. This includes experimenting with different similarity metrics, encoding strategies, and transformations implemented in HPVM-HDC evaluated in Section  5.3. To evaluate the improved programmability of approximations in HDC++, we manually implemented three of the automatic transformations described in Table 3 in the GPU baseline code for HD-Classification. For configuration X, only 4 lines of code needed to be modified. Implementing II required 25 lines of code and approximately 30 minutes of work. Finally, implementing III required the most source code modifications—this configuration requires modifying the allocations of the class hypermatrix and encoded hypervectors, implementing packing elements to 1-bit, and implementating Hamming distance on 1-bit elements. 33 lines of code were written requiring approximately 1 hour. In contrast, HPVM-HDC enables exploring these transformations within seconds by modifying 1-2 lines of the application code. Furthermore, these transformations occur at the HPVM-HDC IR representation, allowing for CPU and GPU implementations of the approximations to be automatically generated, whereas the manual modifications are specific to the GPU baseline.
@@ 6. Related Work
Libraries for HDC. Work on libraries for HDC is relatively extensive. TorchHD (Heddes et al., 2022) is a library implemented using PyTorch (Paszke et al., 2019b) and provides interfaces for implementing HDC algorithms on CPUs and GPUs. OpenHD (Kang et al., 2022b) is a library implemented in Python with CUDA extensions to facilitate implementation of HD Classification and Clustering for GPUs. HDTorch (Simon et al., 2022) is another library implemented in Python with CUDA extensions for hypervector operations for classical and online HDC learning. None of these libraries support nor can be easily extended to target HDC accelerators.
Compilers for HDC. The HDCC compiler (Vergés et al., 2023) is the first and, to date, the only compiler developed for HDC. The primary limitation of HDCC is that it does not support heterogeneous compilation: it only supports one back end that generates C code with POSIX (Walli, 1995) and vector extensions for multi-core CPUs from high-level descriptions of HDC algorithms. Unlike HPVM-HDC, it does not target other hardware architectures, which present more lucrative performance and energy efficiency trade-offs.
Compilers for Deep Learning. Work on compilers for deep learning is extensive (Sabne, 2020; Chen et al., 2018; Rotem et al., 2018; Li et al., 2021). These compilers operate on high-level descriptions of deep learning models, often defined in high-level frameworks such as Tensorflow (Abadi et al., 2016) and PyTorch (Paszke et al., 2019a). Programming HDC algorithms requires tweaking and exploring different algorithmic and approximation choices which can only be facilitated by allowing programmers to express custom algorithms in addition to providing pre-ordained high-level HDC abstractions. This flexibility, however, is not afforded by deep learning frameworks. Moreover, deep learning models are compiled to computational dataflow graphs in XLA (xla, [n. d.]); this representation does not adequately capture all forms of parallelism, unlike HPVM-HDC IR which supports a hierarchical dataflow representation that represents parallelism at different levels of the hierarchy.

### v3

@@ 5.3. Impact of HDC optimizations
For this evaluation, we set the encoding dimension to 10240 and derived the trained class hypervectors using cosine similarity in a single pass over the training dataset offline. Figure 7 plots the relative speed-up versus inference accuracy for different optimization settings described in Table 3. These configurations are not meant to be exhaustive but merely illustrative of points in the various stages in an HDC application which can be approximated. The baseline (I) achieves an inference accuracy of 0.82. While all optimization settings achieve a significant performance improvement, the inference accuracy varies drastically ranging from a low of 0.2 to a high of 0.89. The points in green (III, VII, VIII) result in improved accuracy over the baseline, as well as faster performance. When using the binarization approximation, i.e., projecting the floating point values in the class hypervectors to {+1, -1}, which are represented using 1-bit elements, and using Hamming distance similarity measure, we observe a negligible change in accuracy. This aligns with prior observations regarding the effects of binarized learning on accuracy (Imani et al., 2019c). Additionally, this reduces the initial data movement necessary as the class hypervectors are represented using fewer bits in total. Reduction perforation adds upon the performance improvements achieved from automatic binarization. However, the downstream effect on accuracy varies significantly depending on which operation is being perforated. Configurations (VII, VIII, X) perforate the inference’s similarity computation and result in at most a 13% drop in accuracy with respect to the baseline in X and no drop in the other two. However, when perforating the encoding stages (V, VI, IX), the inference accuracy drops to 25-35%. This implies that for HD classification, the similarity computation is much more amenable to approximation than the encoding stage.
@@ 5.4. Programmability
Lines of Code: The high-level HDC operations in HDC++ also enable more concise code. We compare the lines of code required to implement the applications with their respective baselines. Note that for each target architecture, there is a different baseline version, often implemented in a different programming language. Table  4 describes the total lines of code reduction in HDC++ versus the baselines for the applications. HDC++ provides a 1.6x geomean reduction in total lines of code (by combining the lines of codes across all baseline target implementations) for HDC applications while remaining retargetable. The HD-Classification and HD-Hashtable baseline CPU codes, which are implemented in Python, are smaller than the HDC++ application codes. This is a consequence of HDC++ being embedded in C++ which is inherently more verbose than Python. However, the HD-Classification, HD-Clustering, and HyperOMS baseline GPU codes, which are implemented in CUDA C++, are larger than the HDC++ application codes; HDC++ abstracts away data-movement between the device and host memory. Additionally, the high-level HDC++ primitives compactly represent the corresponding manually written kernel functions in the GPU baselines, resulting in fewer lines of code.
Approximation Configuration: The HDC++ programming interface and HPVM-HDC compiler make it easy to fine-tune HDC applications to balance performance and accuracy because the approximation configuration can be separately specified with almost no changes to the application source code. This includes experimenting with different similarity metrics, encoding strategies, and transformations implemented in HPVM-HDC evaluated in Section  5.3.
To evaluate the improved programmability of approximations in HDC++, we manually implemented three of the automatic transformations described in Table 3 in the GPU baseline code for HD-Classification. For configuration X, only 4 lines of code needed to be modified. Implementing II required 25 lines of code and approximately 30 minutes of work. Finally, implementing III required the most source code modifications—this configuration requires modifying the allocations of the class hypermatrix and encoded hypervectors, implementing packing elements to 1-bit, and implementating Hamming distance on 1-bit elements. 33 lines of code were written requiring approximately 1 hour. In contrast, HPVM-HDC enables exploring these transformations within seconds by modifying 1-2 lines of the application code. Furthermore, these transformations occur in the HPVM-HDC IR representation, allowing for CPU and GPU implementations of the approximations to be automatically generated, whereas the manual modifications are specific to the GPU baseline.
@@ 6. Related Work
Libraries for HDC. Work on libraries for HDC is relatively extensive. TorchHD (Heddes et al., 2022) is a library implemented using PyTorch (Paszke et al., 2019b) and provides interfaces for implementing HDC algorithms on CPUs and GPUs. OpenHD (Kang et al., 2022b) is a library implemented in Python with CUDA extensions to facilitate implementation of HD Classification and Clustering for GPUs. HDTorch (Simon et al., 2022) is another library implemented in Python with CUDA extensions for hypervector operations for classical and online HDC learning. None of these libraries support, nor can be easily extended to target, HDC accelerators.
Compilers for HDC. The HDCC compiler (Vergés et al., 2023) is the first and, to date, the only compiler we know of developed for HDC. However, HDCC supports a very limited programming model. HDCC supports a fixed and limited set of encoding, training, and inference algorithms, rather than a set of lower level hypervector and hypermatrix primitives like HPVM-HDC. Another limitation of HDCC is that it does not support heterogeneous compilation: it only supports one back end that generates C code with POSIX (Walli, 1995) and vector extensions for multi-core CPUs from high-level descriptions of HDC algorithms. Unlike HPVM-HDC, it does not target other hardware architectures, which present more lucrative performance and energy efficiency trade-offs.
Compilers for Deep Learning. Work on compilers for deep learning is extensive (Sabne, 2020; Chen et al., 2018; Rotem et al., 2018; Li et al., 2021). These compilers operate on high-level descriptions of deep learning models, often defined in high-level frameworks such as Tensorflow (Abadi et al., 2016) and PyTorch (Paszke et al., 2019a). Programming HDC algorithms requires tweaking and exploring different algorithmic and approximation choices which can only be facilitated by allowing programmers to express custom algorithms in addition to providing pre-ordained high-level HDC abstractions. This flexibility, however, is not afforded by deep learning frameworks. Moreover, deep learning models are compiled to computational dataflow graphs in XLA (xla, [n. d.]); this representation does not adequately capture all forms of parallelism, unlike HPVM-HDC IR which supports a hierarchical dataflow representation that captures multiple kinds of parallelism.
@@ 7. Conclusion
Acknowledgements.
We thank the anonymous reviewers for their helpful feedback on this paper. This work was supported by funding from PRISM, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.